Font Test Document: pifont

This is an automatically generated document to demonstrate the ‘pifont’ font family. To use this font in your own documents, simply include the following line in the preamble: “\usepackage{pifont}”.

The pifont Alphabet

a b c d e f g h i j k l m n o p q r s t u v w x y z A B C D E F G H I J K L M N O P Q R S T U V W X Y Z

0 1 2 3 4 5 6 7 8 9 ! @ “ ” # $ % ˆ & ‘ ’ ( ) * + , - . / : ; > = < ? \ | \{ \} ˜ £ \begin{matrix}c \\ ◯ \end{matrix}

Sample Text

“But why, some say, the moon? Why choose this as our goal? And they may well ask why climb the highest mountain? Why, 35 years ago, fly the Atlantic? Why does Rice play Texas?

“We choose to go to the moon in this decade and do the other things, not because they are easy, but because they are hard, because that goal will serve to organize and measure the best of our energies and skills, because that challenge is one that we are unwilling to postpone, and one which we intend to win; and the others, too.”

(President John F. Kennedy, 1962 JFK Rice University Speech)

Accented Text

Lowercase: The phœnix’s naïve rôle vis-à-vis soufflés
Uppercase: THE PHŒNIX’S NAÏVE RÔLE VIS-À-VIS SOUFFLÉS

Example Typeface Styles

Plain:  ABCDEFGHIJKLMabcdefghijklm1234567890
Bold:  ABCDEFGHIJKLMabcdefghijklm1234567890
Emphasised:  ABCDEFGHIJKLMabcdefghijklm1234567890
Italic:  ABCDEFGHIJKLMabcdefghijklm1234567890
Bold Empasised: ABCDEFGHIJKLMabcdefghijklm1234567890
Bold Italic:  ABCDEFGHIJKLMabcdefghijklm1234567890

Various Font Sizes

\footnotesize ABCDEFGHIJKLMabcdefghijklm1234567890
\small  ABCDEFGHIJKLMabcdefghijklm1234567890
\normalsize  ABCDEFGHIJKLMabcdefghijklm1234567890
\large  ABCDEFGHIJKLMabcdefghijklm1234567890
\Large  ABCDEFGHIJKLMabcdefghijklm1234567890